I would like some assistance with this problem please.
A metal stays solid at Fahrenheit tempretures below 1817.2 degrees. Determine (in terms of inequality) those Celsius tempretures for which the metal stays solid. Use the formula
F=9/5 C+32.

Answer by stanbon(75887)

A metal stays solid at Fahrenheit tempretures below 1817.2 degrees. Determine (in terms of inequality) those Celsius tempretures for which the metal stays solid. Use the formula
F=9/5 C+32.
------------------------
1817.2 = (9/5)C + 32
(9/5)C = 1785.2
C = 991.7778 degrees Celsius
==================
Cheers,
Stan H.